Given that E° = 0.52V for the reduction Cu^+ (aq) + e^- —> Cu (s), calculate E°, ∆G° and K for the following reaction at 25°C.

2Cu^+ (aq) —> Cu ^2+ (aq) + Cu(s)
1
Expert's answer
2019-09-09T04:53:32-0400

We have the formula:

Eocell=Eocathode−Eoanode Eocell=0.520−0.159 Eocell=0.361V Ecello=Ecathodeo−Eanodeo Ecello=0.520−0.159 Ecello=0.361V

We have the formula,

ΔGo=−nFEo ΔGo=−(1)(96500)(0.361)ΔGo=−34836.5J ΔGo=−nFEo ΔGo=−(1)(96500)(0.361) ΔGo=−34836.5J

Now,

We have,

ΔGo=−2.3022nRTlogK−34836.5=−(2.303)(8.314)(1)(298)logKK=1.27×106